# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ 4.4  Error bounds for the controllable diffusion models
$\quad+\underbrace{\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{base}}^{2}}_{\text{(II)}}+\underbrace{d\exp(-T)}_{\text{(III)}}+\underbrace{\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}}_{\text{(IV)}},$
@@
The proof of Lemma 4.4 is deferred to Appendix D.1. Lemma 4.4 decomposes the KL-divergence between the early-stopping distribution $q_{T_{0}}$ and the distribution of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ into four components: (I) Doob’s guidance error, (II) the base score error, (III) the initialization error, and (IV) the discretization error. Specifically, Doob’s guidance error represents the average error of Doob’s $h$-guidance estimator at each time point, which has been investigated in Theorem 4.3; the base score error is the average error of the base score estimator at each time point, which is discussed in Assumption 3; the initialization error arises from replacing the initial distribution $q_{T}=p_{T}$ with a Gaussian distribution in (3.23); and the discretization error is induced by the exponential integrator.
@@
(4.4) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\Big\{\sigma_{T_{0}}^{2}+\varepsilon+\frac{\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log\Big(\frac{1}{\varepsilon}\Big)\Big\},$
provided that the truncation radius $R$, the step size $h$, the number of steps $K$, the error of base score $\varepsilon_{\mathrm{base}}$, and the number of samples $n$ for Doob’s matching are set, respectively, as
$R=\Theta\Big(\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad h=\Theta\Big(\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad K=\Theta\Big(\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big)\Big),$
$\varepsilon_{\mathrm{base}}^{2}=\Theta\Big(\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon}\Big)\Big),\quad n=\Theta\Big(\frac{1}{\varepsilon^{6(d+4)}}\log^{d+4}\Big(\frac{1}{\varepsilon}\Big)\Big),$
Further, letting $T_{0}=\Theta(\varepsilon)$, it holds that
(4.5) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log\Big(\frac{1}{\varepsilon}\Big).$
@@
The proof of Theorem 4.5 is deferred to Appendix D.2. In this theorem, (4.4) illustrates a trade-off with respect to the early-stopping time $T_{0}$. Specifically, as $T_{0}$ approaches zero, the early-stopping error ${\mathcal{O}}(\sigma_{T_{0}}^{2})$ vanishes, while the second term blows up. By balancing these competing terms, we obtain the optimal convergence rate in (4.5) and provide theoretical guidance for the choice of the early-stopping time. This theorem provides theoretical guidance for the selection of hyper-parameters, including the truncation radius $R$ in (3.24), the step size $h$, the number of steps $K$ in (3.23), the error of base score $\varepsilon_{\mathrm{base}}$ in Assumption 3, and the number of samples $n$ for Doob’s matching in (3.20).
@@ 5  Conclusions
In this work, we proposed Doob’s matching, a principled inference-time alignment framework for diffusion models grounded in the theory of Doob’s $h$-transform. Our approach reformulates guidance as the gradient of the logarithm of an underlying Doob’s $h$-function, providing a mathematically consistent mechanism for tilting a pre-trained diffusion model toward a target distribution without retraining the base score network. By leveraging gradient-penalized regression, Doob’s matching simultaneously estimates both the $h$-function and its gradient, thereby providing a consistent estimator for Doob’s guidance.
From a theoretical perspective, we established non-asymptotic convergence rates for the proposed guidance estimator, showing that the estimated Doob’s guidance converges to the true guidance under suitable choices of the hypothesis class and regularization parameter. Building on this result, we further derived non-asymptotic convergence guarantees for the induced controllable diffusion process, demonstrating that the generated distribution converges to the target distribution in the $2$-Wasserstein distance. These results provide an end-to-end theoretical guarantees for inference-time aligned diffusion models that explicitly account for guidance estimation error, base score estimation error, initialization bias, and discretization error.
@@ Appendix A Derivations in Section 2
Lemma A.1 (Exponential integrator).
Suppose $\mathbf{X}_{t}$ satisfies a linear SDE:
(A.1) $\mathrm{d}\mathbf{X}_{t}=(\mathbf{X}_{t}+\mathbf{b}(s,\mathbf{X}_{s}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},\quad t\geq s,$
Then it follows that
$\mathbf{X}_{t}=\exp(t-s)\mathbf{X}_{s}+\phi^{2}(t-s)\mathbf{b}(s,\mathbf{X}_{s})+\phi(2(t-s))\bm{\xi},$
where $\bm{\xi}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ and $\phi(z)=\sqrt{\exp(z)-1}$.
Appendix B Derivations in Section 3
Lemma B.1.
@@ Appendix B Derivations in Section 3
(B.1) $h^{*}(t,\mathbf{x})\coloneq{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})\mid\mathbf{X}_{t}^{\leftarrow}=\mathbf{x}].$
@@
(B.2) $=L_{t}(\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},$
where the second equality is due to Lemma B.1. Using Itô’s formula for log-likelihood yields
@@
where the second equality is due to (B.2). This completes the proof. ∎
@@
Let the base process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
(B.3) ${\mathbb{E}}^{{\mathbb{P}}}\Big[\exp\Big(\int_{0}^{T}\|\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}s\Big)\Big]<\infty.$
@@
(B.4) $\mathrm{d}\widetilde{\mathbf{B}}_{t}=\,\mathrm{d}\mathbf{B}_{t}-\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t,$
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the base process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
(B.5) $\mathrm{d}\mathbf{X}_{t}^{\leftarrow}=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})+2\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\widetilde{\mathbf{B}}_{t}.$
@@
Step 2. Dynamics under the target path measure. Recall that under the base path measure $\mathbb{P}$}, the process evolves as

### v2

@@ 5.3  Error bounds for the controllable diffusion models
$\quad+\underbrace{\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{ref}}^{2}}_{\text{(II)}}+\underbrace{d\exp(-T)}_{\text{(III)}}+\underbrace{\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}}_{\text{(IV)}},$
@@
The proof of Lemma 5.5 is deferred to Appendix E. Lemma 5.5 decomposes the KL-divergence between the early-stopping distribution $q_{T_{0}}$ and the distribution of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ into four components: (I) Doob’s guidance error, (II) the reference score error, (III) the initialization error, and (IV) the discretization error. Specifically, Doob’s guidance error represents the average error of Doob’s $h$-guidance estimator at each time point, which has been investigated in Theorem 5.3; the reference score error is the average error of the reference score estimator at each time point, which is discussed in Assumption 3; the initialization error arises from replacing the initial distribution $q_{T}=p_{T}$ with a Gaussian distribution in (4.8); and the discretization error is induced by the exponential integrator.
@@
(5.3) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big),$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
@@
The proof of Theorem 5.6 is deferred to Appendix E. This theorem establishes a non-asymptotic convergence rate for the controllable diffusion model (4.8) using the variationally stable Doob’s matching (4.6). Crucially, it provides theoretical guidance for selecting hyper-parameters, including the truncation radius $R$ in (4.9), the step size $h$, the number of steps $K$ in (4.8), the early stopping time $T_{0}$, the terminal time $T$, the reference score error $\varepsilon_{\mathrm{ref}}$ (Assumption 3), and the sample size $n$ for Doob’s matching in (4.4).
However, this rate suffers from the curse of dimensionality (CoD), implying that the required number of samples $n$ grows exponentially as the error tolerance $\varepsilon$ decays. We address this challenge in the remainder of this section under a low-dimensional subspace assumption.
5.4  Adaptivity to low-dimensionality
In this subsection, we demonstrate that the convergence rate mitigates the curse of dimensionality under a low-dimensional subspace assumption, a setting previously explored in Chen et al. (2023b, Section 3) and Oko et al. (2023, Section 6).
Assumption 4 (Low-dimensional subspace).
Let $d^{*}\ll d$ be an integer, and $\mathbf{P}\in{\mathbb{R}}^{d\times d^{*}}$ be a column orthogonal matrix, i.e., $\mathbf{P}^{\top}\mathbf{P}=\mathbf{I}_{d^{*}}$. Let $\bar{p}_{0}$ be a probability density with a compact support contained within a hypercube $\{\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d^{*}}:\|\bar{\mathbf{x}}_{0}\|_{\infty}\leq 1\}$. The reference density $p_{0}$ is a push-forward of $\bar{p}_{0}$ by the linear map $\mathbf{P}$, i.e., $p_{0}\coloneq\mathbf{P}_{\sharp}\bar{p}_{0}$.
Consequently, the reference density $p_{0}$ is supported on a linear subspace $\{\mathbf{P}\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d}:\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d^{*}}\}$ with an ambient dimension $d$, and a much smaller intrinsic dimension $d^{*}\ll d$.
Before proceeding, we define the forward and time-reversal process in the low-dimensional latent space. Analogously to (2.1), the forward process reads
(5.4) $\mathrm{d}\bar{\mathbf{X}}_{t}=-\bar{\mathbf{X}}_{t}\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\bar{\mathbf{B}}_{t},\quad t\in(0,T),\penalty 10000\ \bar{\mathbf{X}}_{0}\sim\bar{p}_{0},$
where $\bar{\mathbf{B}}_{t}$ is a $d^{*}$-dimensional standard Brownian motion, and $T>0$ is the terminal time. The transition distribution of this forward process is given by:
(5.5) $(\bar{\mathbf{X}}_{t}|\bar{\mathbf{X}}_{0}=\bar{\mathbf{x}}_{0})\sim\mathcal{N}(\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d^{*}}).$
Let $\bar{p}_{t}$ denote the marginal density of $\bar{\mathbf{X}}_{t}$ for $t\in(0,T)$. The corresponding time-reversal process (Anderson, 1982) is defined as
(5.6) $\,\mathrm{d}\bar{\mathbf{X}}_{t}^{\leftarrow}$ $=(\bar{\mathbf{X}}_{t}^{\leftarrow}+2\nabla\log\bar{p}_{T-t}(\bar{\mathbf{X}}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\bar{\mathbf{B}}_{t},\quad t\in(0,T),$
$\bar{\mathbf{X}}_{0}^{\leftarrow}$ $\sim\bar{p}_{T}.$
As shown by Anderson (1982), the path measure of the time-reversal process $(\bar{\mathbf{X}}_{t}^{\leftarrow})_{0\leq t\leq T}$ corresponds exactly to the reverse of the forward process $(\bar{\mathbf{X}}_{t})_{0\leq t\leq T}$.
The following result establishes a relationship between the ground-truth Doob’s $h$-function (3.10) and its analogue $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\to{\mathbb{R}}$ in the low-dimensional latent space. In other words, it provides a low-dimensional representation of the ground-truth Doob’s $h$-function.
Proposition 5.7 (Low-dimensional representation).
Suppose Assumptions 4 and 2 hold. Then for any $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, we have
(5.7) $h_{t}^{*}(\mathbf{x})=\bar{h}_{t}^{*}(\mathbf{P}^{\top}\mathbf{x})\coloneq{\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}].$
The proof of Proposition 5.7 is provided in Appendix F. Proposition 5.7 implies that estimating the ground-truth Doob’s $h$-function reduces to estimating its low-dimensional counterpart $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\rightarrow{\mathbb{R}}$, thereby enabling the Doob’s guidance estimator to adapt to low-dimensional structures.
Analogously to Proposition 5.1, we can establish the regularity properties of the low-dimensional representation of Doob’s $h$-function $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\to{\mathbb{R}}$ in (5.7).
Proposition 5.8.
Suppose Assumptions 4 and 2 hold. Then for all $t\in(0,T)$ and $\bar{\mathbf{x}}\in{\mathbb{R}}^{d^{*}}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq\bar{h}_{t}^{*}(\bar{\mathbf{x}})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
The proof of Proposition 5.7 is provided in Appendix F. Based on these results, we derive the convergence rates for the variationally stable Doob’s matching under the assumption of low-dimensional subspace.
Theorem 5.9 (Adaptivity to intrinsic dimension).
(5.8) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d^{*}+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d^{*}+8}})$, where $C$ is a constant depending only on $d^{*}$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 5.9 is provided in Appendix F. This result confirms that the convergence rate eliminates the exponential dependence on the ambient dimension $d$, scaling exponentially solely with the intrinsic dimension $d^{*}\ll d$. This effectively mitigates the curse of dimensionality in Theorem 5.3.
The following corollary is a direct consequence of Theorem 5.9, derived using arguments similar to those in Theorem 5.6.
Corollary 5.10.
(5.9) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big).$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d^{*}+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
Here $C$ is a constant depending only on $^{*}, dd$, $\bar{B}$, and $\underline{B}$.
Crucially, the convergence rates in Corollary 5.10 depend only polynomially on the ambient dimension $d$, while the sample complexity depends exponentially solely on the intrinsic dimension $d^{*}\ll d$. This result significantly mitigates the curse of dimensionality.
Remark 5.11.
Adaptivity to low dimensionality plays a pivotal role in the analysis of diffusion and flow-based models. One line of work studies the adaptivity of score or velocity estimator to low dimensionality under Assumption 4 or its variants; see e.g., Chen et al. (2023b); Oko et al. (2023); Yakovlev and Puchkin (2025a); Ding et al. (2025a). A second line of work focuses on the adaptivity of the sampling procedure; see e.g., Li and Yan (2024); Huang et al. (2024); Potaptchik et al. (2025). These works provide valuable insights for extending Corollary 5.10 to achieve provable adaptivity across reference score estimation, sampling, and guidance estimation, thereby completely eliminating dependence on the ambient dimension. We leave this unified analysis, which is outside the scope of the current work, for future research. Importantly, even within the current framework, the dependence on the ambient dimension $d$ remains only polynomial.
@@ 6  Conclusions
In this work, we proposed variationally stable Doob’s matching, a principled inference-time alignment framework for diffusion models grounded in the theory of Doob’s $h$-transform. Our approach reformulates guidance as the gradient of the logarithm of an underlying Doob’s $h$-function, providing a mathematically consistent mechanism for tilting a pre-trained diffusion model toward a target distribution without retraining the reference score network. By leveraging gradient-regularized regression, Doob’s matching simultaneously estimates both the $h$-function and its gradient, thereby providing a consistent estimator for Doob’s guidance.
From a theoretical perspective, we established non-asymptotic convergence rates for the proposed guidance estimator, showing that the estimated Doob’s guidance converges to the true guidance under suitable choices of the hypothesis class and regularization parameter. Building on this result, we further derived non-asymptotic convergence guarantees for the induced controllable diffusion process, demonstrating that the generated distribution converges to the target distribution in the $2$-Wasserstein distance. These results provide an end-to-end theoretical guarantees for inference-time aligned diffusion models that explicitly account for guidance estimation error, reference score estimation error, initialization bias, and discretization error. Furthermore, we show that our convergence rates depend solely on the intrinsic dimension of the linear subspace rather than the ambient dimension. This highlights the estimator’s adaptivity to low-dimensional structures, effectively mitigating the curse of dimensionality.
@@ Appendix A Derivations in Section 3
Lemma A.1.
@@
(A.1) $h^{*}(t,\mathbf{x})\coloneq{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})\mid\mathbf{X}_{t}^{\leftarrow}=\mathbf{x}].$
@@
(A.2) $=L_{t}(\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},$
where the second equality is due to Lemma A.1. Using Itô’s formula for log-likelihood yields
@@
where the second equality is due to (A.2). This completes the proof. ∎
@@
Let the reference process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
(A.3) ${\mathbb{E}}^{{\mathbb{P}}}\Big[\exp\Big(\int_{0}^{T}\|\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}s\Big)\Big]<\infty.$
@@
(A.4) $\mathrm{d}\widetilde{\mathbf{B}}_{t}=\,\mathrm{d}\mathbf{B}_{t}-\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t,$
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the reference process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
(A.5) $\mathrm{d}\mathbf{X}_{t}^{\leftarrow}=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})+2\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\widetilde{\mathbf{B}}_{t}.$
@@
Step 2. Dynamics under the target path measure. Recall that under the reference path measure $\mathbb{P}$}, the process evolves as